peter wanted to draw a line parallel to the line of a window with equation y=x+2.if the line is supposed to pass through point (2,3).what is the equation of the line desired by peter.

Answer by venugopalramana(3286)

y=x+2....line parallel to this is given by y=x+k..it is passing through (2,3)
so substituting...3=2+k..or k=3-2=1
hence the eqn.of the parallel line is y=x+1